# Markovian dynamics for a quantum/classical system and quantum trajectories

### v1

@@ 4.2 Jumps of discrete type
$\mathcal{K}[a\otimes f](x)=f(x)\widetilde{\mathcal{L}}(x)[a]+a\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+a\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}a_{i}^{k}(x)\left(aL_{k}(x)+L_{k}(x)^{\dagger}a\right)+\sum_{k=1}^{l}\lambda_{k}f\big{(}x+g(x,k)\big{)}J_{k}(x)^{\dagger}aJ_{k}(x),$ (50)
@@ 4.2.1 Unitary jumps
$J_{k}(x)=\beta_{k}(x)U_{k}(x),\quad\beta_{k}(x)>0,\quad U_{k}(x)^{\dagger}U_{k}(x)=\mathds{1},\quad\Rightarrow\quad J_{k}(x)^{\dagger}J_{k}(x)=\beta_{k}(x)^{2}\mathds{1},\quad I_{k}(x)=\beta_{k}(x)^{2}.$
@@
Note that $I_{kt}$ does not depend on the quantum state and this gives that, under $\mathbb{P}$, the full law of the processes $N_{k}(t)$ remains independent of the quantum state.
@@ 4.3 Entanglement creation and preservation
To quantify the entanglement we consider the concurrence [55], whose definition is recalled in Appendix B. Now we have the a priori state and the conditional state, which is random. Due to the definition (65) of the concurrence of a mixed state, it is easy to see that the concurrence of the a priori state $\eta_{t}$ is a lower bound for the mean concurrence of the conditional state $\widehat{\psi}_{t}$:
@@
By the SDE (53), we get the evolution equation for the random chi-quantity:
@@
$H=\frac{\omega}{2}\left(\sigma^{1}_{z}+\sigma^{2}_{z}\right),\qquad U_{k}(x)=U_{k}=\begin{cases}\mathrm{e}^{\mathrm{i}\phi_{k}\vec{\sigma}\cdot\vec{\epsilon}^{k}}\otimes\mathds{1}\quad{}&k=1,\ldots,l_{1},\\ \mathds{1}\otimes\mathrm{e}^{\mathrm{i}\phi_{k}\vec{\sigma}\cdot\vec{\epsilon}^{k}}&k=l_{1}+1,\dots,l,\end{cases}\quad\phi_{k}\in\mathbb{R},\quad\vec{\epsilon}^{k}\in\mathbb{R}^{3},\quad\left|\vec{\epsilon}^{k}\right|=1.$
By (54), (66) we get $\mathrm{d}\chi(t)=0$: the conditional concurrence is constant in time. If the initial state is maximally entangled, we have $\operatorname{\mathbb{E}}_{\mathbb{P}}\big{[}C(\widehat{\psi}_{t})\big{]}=1$, $\forall t\geq 0$.
@@
$\omega>0,\qquad l=2,l_{1}=1,\quad\lambda_{k}=\lambda>0,\quad\beta_{k}=\beta\neq 0,\quad\phi_{k}=\frac{\pi}{2},\quad\vec{\epsilon}_{k}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}\qquad\Rightarrow\qquad U_{k}=\mathrm{i}\sigma_{x}^{k}.$
@@
For the two model above we can speak of the phenomenon of hidden entanglement, a term which is sometimes used for situations in which strong entanglement is present in a statistical ensemble, while it is less relevant in the associated mean state, see for instance [42, 29].
@@ 4.4 Control
Moreover, we have that the classical subsystem satisfies Eq. (2.3.1) under the physical probability $\mathbb{P}$; so, its dynamics depends on the state of the quantum subsystem through the coefficients $C^{(i)}(t)$. On the other side, the classical component can affect the quantum dynamics through the $x$-dependence of the operators $H(x)$, $L_{k}(x)$, $J(x,u)$. In this way we have realized a closed loop feedback. This opens the question of the connections with quantum control theory, where already the quantum trajectory formalism (quantum filtering) has been explored, see [56] and references therein.
@@ 5 Conclusions
In the present approach, the hybrid dynamics has been introduced by a couple of SDEs, one for the classical component and one for the quantum component, and this has been done in two different ways. In the first case a reference probability $\mathbb{Q}$ is used, and the two SDEs are the evolution equation (2) for the classical process and the linear SME (12) for the (non-normalized) quantum state; the quantum/classical interaction is contained only in the SME (12). The physical probability $\mathbb{P}$ is determined by the dynamics itself, and it is defined in (14a), (15). By suitable hypothesis (Assumptions 1, 2, 3) it is possible to prove existence and uniqueness of the solution of these coupled equations.
An equivalent description can be given by working under the physical probability $\mathbb{P}$; now we have the SDE (18a) for the classical process and the non-linear SME (19) for the quantum states. However, the mathematical setting is more involved as even the probability law depends on the solution of the two coupled equations. On the other side, in this formulation the quantum/classical interaction is more transparent, as it appears in both equations.
We have also shown that the general axiomatic and the linear structure of a quantum theory is respected. Indeed, we have shown how to connect the whole construction to the general notion of quantum observables (POVMs) and to the notion of “instrument” (Sec. 2.4.2). We have shown that, as expected, the extraction of information from the quantum subsystem is necessarily connected to a suitable dissipation in the dynamics. The absence of memory is underlined by the fact that probabilities and dynamics can be summarized in a notion of “transition instruments” (26), a quantum analogue of the transition probabilities in the theory of classical Markov processes.
Moreover, the Markovian character of the dynamics appears also in the fact that it can be expressed as a hybrid dynamical semigroup. However, in the pure classical case it is shown that the classical states can not always be reduced to densities with respect to the Lebesgue measure (Appendix A); then, the hybrid semigroup needs a $C^{*}$-algebraic formulation. In any case, we have given its formal generator, which allows to compare our results with less general approaches based on master equations.
@@ Appendix A The classical Markov process under the reference probability $\mathbb{Q}$
In this appendix we collect known results on the Markov process defined by (2), when there is no interaction with the quantum component; in particular, we discuss the associated semigroup. Equation (2) is studied in [39, Sec. 6.2].
By following [39, Secs. 6.4.1, 6.4.2], we can introduce the stochastic flow $\Phi$ associated to the SDE (2). For $0\leq s\leq t$, let $\Phi^{s}_{t}(x)$ be the solution of (2) with initial condition $x$ at time $s$; in particular $\Phi^{0}_{t}(x)=X^{x}(t)$. The solution of (2) is an homogeneous Markov process and we have
@@
Then, the transition probabilities (5) satisfy the properties of a time homogeneous transition function [38, Chapt. 4 Sec. 1 p. 156]:  (a) $P(t,x,\cdot)$ is a probability measure on $\mathcal{B}(\mathbb{R}^{s})$, $\forall t\in[0,+\infty)$, $\forall x\in\mathbb{R}^{s}$;  (b) $P(0,x,\cdot)=\delta_{x}(\cdot)$, $\forall x\in\mathbb{R}^{s}$;  (c) $P(\cdot,\cdot,\mathsf{E})$ is a bounded Borel function on $[0,+\infty)\times\mathbb{R}^{s}$, $\forall\mathsf{E}\in\mathcal{B}(\mathbb{R}^{s})$;  (d) the Chapman-Kolmogorov identity holds, which is
@@
$\{T_{t},\,t\geq 0\}$ is a semigroup of positive, bounded linear maps from $B_{\rm b}(\mathbb{R}^{s})$ into itself, with $T_{t}(1)=1$, $T_{0}=\mathrm{Id}$. Moreover, $T_{t}$ maps $C_{\rm b}(\mathbb{R}^{s})$ into itself [39, Note 3 p. 402]. The expression (58) is continuous in $t$, for fixed $f$ and $x$.
@@
$\mathrm{d}f\big{(}X(t)\big{)}\Big{|}_{X(t_{-})=x}=\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\left(c_{i}(x)\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}(x)\mathrm{d}W_{k}(t)\right)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\mathrm{d}t\\ {}+\mathrm{d}t\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\\ {}+\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\Pi(\mathrm{d}u,\mathrm{d}t).$ (59)
@@
$\mathcal{K}^{\rm cl}[f](x)=\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u).$ (61)
Under some additional sufficient assumptions, it is possible to prove that $T_{t}$ is strongly continuous and that it is generated by $\mathcal{K}^{\rm cl}$, cf. [39, (6.36) p. 402]. We do not need these results, because existence and uniqueness of the solution of the SDE (2) is enough to give the intrinsic dynamics of the classical component of the hybrid system. In [37, pp. 56, 349-350], (60) is used as definition of “generator of the time homogeneous Markov process”; the law of the process is determined by the Dynkin’s expectation formula
@@
In some cases the adjoint of $T_{t}$, defined in the dual Banach space, leaves invariant the space of the probability densities with respect to the Lebesgue measure and some known classical master equations are obtained. In the deterministic case, i.e. $a_{i}^{k}(x)=0$ and $g(x,u)=0$, we can take $s$ even and the coefficients $c_{i}(x)$ giving an Hamiltonian evolution; then, the master equation is the Liouville equation. Instead, with only $g(x,u)=0$, we can get the Kolmogorov-Fokker-Planck equation [39, Sec. 3.5.3], [36, Secs. 9.3, 14.2.2].
In general, it can happen that $X(t)$ has not a density with respect to Lebesgue measure, even if we start with $X(0)$ having a density. As an example, take the equation $\mathrm{d}X(t)=\left[x^{0}-X(t_{-})\right]\mathrm{d}N(t)$ with $x^{0}\in\mathbb{R}^{s}$; $X(0)$ is taken with a density and independent of the Poisson process $N$ of intensity $\lambda>0$. Then, we obtain $P[X(t)=x^{0}]=\left(1-\mathrm{e}^{-\lambda t}\right)$ and a density with respect to Lebesgue does not exist for $t>0$.
By generalizing the example above we can construct models with jumps among fixed points at random times. Let us take $n$ distinct points $x^{j}\in\mathbb{R}^{s}$ and $l$ independent Poisson processes $N_{k}$ of intensities $\lambda_{k}>0$,
@@
After the first jump, the process reduces to a random walk on the points $\{x^{1},\ldots,x^{n}\}$. When the initial condition has a probability concentrated in the set of points $\{x^{1},\ldots,x^{n}\}$, the probability at time $t$ satisfies the Pauli master equation [2, (13)].

### v2

@@ 4.2 Jumps of discrete type
$\mathcal{K}[a\otimes f](x)=f(x)\widetilde{\mathcal{L}}(x)[a]+a\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+a\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}b_{i}^{k}(x)\left(aL_{k}(x)+L_{k}(x)^{\dagger}a\right)+\sum_{k=1}^{l}\lambda_{k}f\big{(}x+g(x,k)\big{)}J_{k}(x)^{\dagger}aJ_{k}(x),$ (57)
@@
$J_{k}(x)=\beta_{k}(x)U_{k}(x),\quad\beta_{k}(x)>0,\quad U_{k}(x)^{\dagger}U_{k}(x)=\mathds{1},$
$\Rightarrow\quad J_{k}(x)^{\dagger}J_{k}(x)=\beta_{k}(x)^{2}\mathds{1},\quad I_{k}(x)=\beta_{k}(x)^{2}.$
@@
Note that $I_{kt}=\beta_{k}\big{(}X(t_{-})\big{)}^{2}$ does not depend on the quantum state and this gives that, under $\mathbb{P}$, the full law of the processes $N_{k}(t)$ remains independent of the quantum state.
@@ 4.3 Entanglement creation and preservation
To quantify the entanglement we consider the concurrence [62], whose definition is recalled in B. Now we have the a priori state and the conditional state, which is random. Due to the definition (72) of the concurrence of a mixed state, it is easy to see that the concurrence of the a priori state $\eta_{t}$ is a lower bound for the mean concurrence of the conditional state $\widehat{\psi}_{t}$:
@@
By (60), we get the evolution equation for the random chi-quantity:
@@
$H=\frac{\omega}{2}\left(\sigma^{1}_{z}+\sigma^{2}_{z}\right),\qquad U_{k}(x)=U_{k}=\begin{cases}\mathrm{e}^{\mathrm{i}\phi_{k}\vec{\sigma}\cdot\vec{\epsilon}^{k}}\otimes\mathds{1}\quad{}&k=1,\ldots,l_{1},\\ \mathds{1}\otimes\mathrm{e}^{\mathrm{i}\phi_{k}\vec{\sigma}\cdot\vec{\epsilon}^{k}}&k=l_{1}+1,\dots,l,\end{cases}$
with $\phi_{k}\in\mathbb{R}$,  $\vec{\epsilon}^{k}\in\mathbb{R}^{3}$,  $\left|\vec{\epsilon}^{k}\right|=1$.  By (61), (73) we get $\mathrm{d}\chi(t)=0$: the conditional concurrence is constant in time. If the initial state is maximally entangled, we have $\operatorname{\mathbb{E}}_{\mathbb{P}}\big{[}C(\widehat{\psi}_{t})\big{]}=1$, $\forall t\geq 0$.
@@
[…]
@@
For the two models above we can speak of the phenomenon of hidden entanglement, a term which is sometimes used for situations in which strong entanglement is present in a statistical ensemble, while it is less relevant in the associated mean state, see for instance [47, 34].
@@ 4.4 Control
Moreover, we have that the classical subsystem satisfies Eq. (20) under the physical probability $\mathbb{P}$; so, its dynamics depends on the state of the quantum subsystem through the coefficients $C^{(i)}(t)$. On the other side, the classical component can affect the quantum dynamics through the $x$-dependence of the operators $H(x)$, $L_{k}(x)$, $J(x,u)$. In this way we have realized a closed loop feedback; applications in quantum optics can be found in [23, 24, 58]. This opens the question of the connections with a full quantum control theory, where already the quantum trajectory formalism (quantum filtering) started to be explored, see [63, 64, 28] and references therein.
@@ 5 Conclusions
In the present approach, the hybrid dynamics has been introduced by a couple of stochastic differential equations, one for the classical component and one for the quantum component, and this has been done in two different ways. In the first case a reference probability $\mathbb{Q}$ is used, and the two stochastic equations are the evolution equation (2) for the classical process and the linear SME (10) for the (non-normalized) quantum state; the quantum/classical interaction is contained only in the SME (10). The physical probability $\mathbb{P}$ is determined by the dynamics itself, and it is defined in (13), (17). By suitable hypothesis (Assumptions 1, 2, 3) it is possible to prove existence and uniqueness of the solution of these coupled equations.
An equivalent description can be given by working under the physical probability $\mathbb{P}$; now we have the stochastic equation (20) for the classical process and the non-linear SME (22) for the quantum states. However, the mathematical setting is more involved as even the probability law depends on the solution of the two coupled equations. On the other side, in this formulation the quantum/classical interaction is more transparent, as it appears in both equations.
We have also shown that the general axiomatic and the linear structure of a quantum theory is respected. Indeed, we have shown how to connect the whole construction to the general notion of quantum observable (the positive operator valued measures) and to the notion of “instrument” (Sec. 2.4.2). We have shown that, as expected, the extraction of information from the quantum subsystem is necessarily connected to the presence of suitable dissipation in the dynamics (see the comments in Sec. 2.3.1, after Eq. (21)). The absence of memory is underlined by the fact that probabilities and dynamics can be summarized in a notion of “transition instruments” (29), a quantum analogue of the transition probabilities in the theory of classical Markov processes.
Moreover, the Markovian character of the dynamics appears also in the fact that it can be expressed as a hybrid dynamical semigroup. However, in the pure classical case it is shown that the classical states can not always be reduced to densities with respect to the Lebesgue measure (A); then, the hybrid semigroup needs a $C^{*}$-algebraic formulation. In any case, we have given its formal generator, which allows to compare our results with less general approaches based on master equations.
@@ Appendix A The classical Markov process under the reference probability $\mathbb{Q}$
In this appendix we collect known results on the Markov process defined by (2), when there is no interaction with the quantum component; in particular, we discuss the associated semigroup. Equation (2) is studied in [44, Sec. 6.2].
By following [44, Secs. 6.4.1, 6.4.2], we can introduce the stochastic flow $\Phi$ associated to the stochastic differential equation (2). For $0\leq s\leq t$, let $\Phi^{s}_{t}(x)$ be the solution of (2) with initial condition $x$ at time $s$; in particular, with the notation introduced at the end of Proposition 1, we have $\Phi^{0}_{t}(x)=X^{x}(t)$. The solution of (2) is an homogeneous Markov process and we have
@@
Then, the transition probabilities (3) satisfy the properties of a time-homogeneous transition function [43, Chapt. 4 Sec. 1 p. 156]:  (a) $P(t,x,\cdot)$ is a probability measure on $\mathcal{B}(\mathbb{R}^{s})$, $\forall t\in[0,+\infty)$, $\forall x\in\mathbb{R}^{s}$;  (b) $P(0,x,\cdot)=\delta_{x}(\cdot)$, $\forall x\in\mathbb{R}^{s}$;  (c) $P(\cdot,\cdot,\mathsf{E})$ is a bounded Borel function on $[0,+\infty)\times\mathbb{R}^{s}$, $\forall\mathsf{E}\in\mathcal{B}(\mathbb{R}^{s})$;  (d) the Chapman-Kolmogorov identity holds, which is
@@
$\{T_{t},\,t\geq 0\}$ is a semigroup of positive, bounded linear maps from $B_{\rm b}(\mathbb{R}^{s})$ into itself, with $T_{t}(1)=1$, $T_{0}=\mathrm{Id}$. Moreover, $T_{t}$ maps $C_{\rm b}(\mathbb{R}^{s})$ into itself [44, Note 3 p. 402]. The expression (65) is continuous in $t$, for fixed $f$ and $x$.
@@
$\mathrm{d}f\big{(}X(t)\big{)}\Big{|}_{X(t_{-})=x}=\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\left(c_{i}(x)\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}(x)\mathrm{d}W_{k}(t)\right)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\mathrm{d}t\\ {}+\mathrm{d}t\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\\ {}+\int_{\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\Pi(\mathrm{d}u,\mathrm{d}t).$ (66)
@@
$\mathcal{K}^{\rm cl}[f](x)=\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,c_{i}(x)+\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\\ {}+\int_{\mathscr{U}}\left(f\big{(}x+g(x,u)\big{)}-f(x)-\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u).$ (68)
Under some additional sufficient assumptions, it is possible to prove that $T_{t}$ is strongly continuous and that it is generated by $\mathcal{K}^{\rm cl}$, cf. [44, (6.36) p. 402]. We do not need these results, because existence and uniqueness of the solution of (2) is enough to give the intrinsic dynamics of the classical component of the hybrid system. In [42, pp. 56, 349-350], (67) is used as definition of “generator of the time-homogeneous Markov process”; the law of the process is determined by the Dynkin’s expectation formula
@@
In some cases the adjoint of $T_{t}$, defined in the dual Banach space, leaves invariant the space of the probability densities with respect to the Lebesgue measure and some known classical master equations are obtained. In the deterministic case, i.e. $b_{i}^{k}(x)=0$ and $g(x,u)=0$, we can take $s$ even and the coefficients $c_{i}(x)$ giving an Hamiltonian evolution; then, the master equation is the Liouville equation. Instead, with only $g(x,u)=0$, we can get the Kolmogorov-Fokker-Planck equation [44, Sec. 3.5.3], [41, Secs. 9.3, 14.2.2].
In general, it can happen that $X(t)$ has not a density with respect to Lebesgue measure, even if we start with $X(0)$ having a density. As an example, take the equation $\mathrm{d}X(t)=\left[x^{0}-X(t_{-})\right]\mathrm{d}N(t)$ with $x^{0}\in\mathbb{R}^{s}$; $X(0)$ is taken with a density and independent of the Poisson process $N$ of intensity $\lambda>0$. Then, we obtain $P[X(t)=x^{0}]=1-\mathrm{e}^{-\lambda t}$ and a density with respect to the Lebesgue measure does not exist for $t>0$.
By generalizing the example above, we can construct models with jumps among fixed points at random times. Let us take $n$ distinct points $x^{j}\in\mathbb{R}^{s}$ and $l$ independent Poisson processes $N_{k}$ of intensities $\lambda_{k}>0$,
@@
After the first jump, the process reduces to a random walk on the points $\{x^{1},\ldots,x^{n}\}$. When the initial condition has a probability concentrated in the set of points $\{x^{1},\ldots,x^{n}\}$, the probability at time $t$ satisfies a Pauli rate equation [2, (13)], [3, (14)].